# Constraining primordial black hole abundance with Insight-HXMT

### v2

@@ Abstract
Primordial black holes (PBHs) are a major candidate for dark matter and they have been extensively constrained across most mass ranges. However, PBHs in the mass range of $10^{17}$$10^{21}$ g remain a viable explanation for all dark matter. In this work, we use observational data from the Hard X-ray Modulation Telescope (Insight-HXMT) to refine constraints on PBHs within the mass range of $2\times 10^{16}$$5\times 10^{17}$ g. Our analysis explores three scenarios: directly using observational data, incorporating the astrophysical background model (ABM), and employing the power-law spectrum with an exponential cutoff. Our results indicate that although Insight-HXMT does not have an advantage in the first two scenarios, when considering the power-law model, its exceptional sensitivity in the hard X-ray regime and sufficiently high upper energy limit significantly strengthen the constraints on PBHs with masses greater than $10^{17}$ g compared to previous limits. Furthermore, the exclusion limit for PBHs as dark matter has reached $4\times 10^{17}$ g, which is comparable to the current threshold.
1 Introduction
@@
Dark matter remains one of the most profound mysteries in modern astrophysics and cosmology. Observations such as galactic rotation curves, gravitational lensing, and anisotropies in the cosmic microwave background provide compelling indirect evidence for its existence. However, despite these clues, the fundamental nature of dark matter continues to elude direct detection. Current models suggest that dark matter constitutes approximately 85% of the total matter in the universe [1, 2], interacting with baryonic matter primarily through gravity and remaining invisible to electromagnetic observations. Over the past few decades, numerous candidates have been proposed, such as weakly interacting massive particles, sterile neutrinos, and axions. These candidates often require modifications to the standard model of particle physics, and none have been definitively detected to date.
@@
Amid these ongoing searches, primordial black holes (PBHs) have emerged as an intriguing alternative, potentially bypassing the need for such modifications. PBHs are black holes that formed in the early universe due to the collapse of high-density fluctuations shortly after the Big Bang [3]. Unlike stellar black holes, PBHs can span a wide range of masses, from the Planck scale to billions of solar masses. This broad mass spectrum and early formation make PBHs relevant not only to the dark matter problem but also to various astrophysical phenomena. For instance, PBHs could account for the unexpected massive early galaxies observed by the James Webb Space Telescope (JWST) [4, 5], and they offer a plausible explanation for the origin of the stellar-mass binary black holes detected by the LIGO and Virgo collaborations [6, 7].
@@
While many mass ranges for PBHs as the entirety of dark matter have been ruled out through observations like lensing, gravitational waves, and accretion, a relatively unconstrained mass window remains between $10^{17}$ g and $10^{21}$[8, 9]. This mass range is particularly notable because it circumvents many of the astrophysical and cosmological constraints that apply to PBHs at both lower and higher masses. In this regime, Hawking radiation becomes a key mechanism for detecting PBHs and constraining their abundance. PBHs in this window are expected to emit radiation primarily in the keV – MeV energy range, making instruments sensitive to hard X-rays and soft $\gamma$-rays ideal for imposing tighter constraints on their abundance.
@@
China’s first X-ray astronomy satellite, the Hard X-ray Modulation Telescope (Insight-HXMT) offers a broad energy range, large effective area, and high sensitivity, particularly in the hard X-ray band [10, 11]. Its sensitivity surpasses that of other instruments, such as the International Gamma-Ray Astrophysics Laboratory/Imager on Board the INTEGRAL Satellite (INTEGRAL/IBIS) and Rossi X-ray Timing Explorer/High-Energy X-ray Timing Experiment (RXTE/HEXTE) [11]. These features make Insight-HXMT particularly well-suited for detecting Hawking radiation from PBHs in the unconstrained mass window, positioning it as an ideal tool for constraining PBH abundance within this range.
@@
In this study, we use diffuse X-ray background data to constrain the abundance of PBHs in the mass range of $2\times 10^{16}$$5\times 10^{17}$ g. Our analysis takes into account both the photon flux generated directly by Hawking radiation and the photons produced from the annihilation of positrons emitted during PBH evaporation. We include contributions to the diffuse X-ray background from both Galactic and extragalactic sources. The data used in our analysis are sourced from the diffuse X-ray background observed by Insight-HXMT [12]. We evaluate three scenarios: directly using the observational data, incorporating an astrophysical background model (ABM), and applying the power-law model fitted from the data to extrapolate the upper energy limit. By leveraging Insight-HXMT’s advanced observational capabilities, our goal is to significantly improve existing constraints on PBH abundance and advance the limits set by prior studies.
@@
The structure of the paper is as follows: In Section 2, we detail the calculations of photon flux from PBH evaporation. Section 3 presents the statistical analysis and constraints on PBH abundance using observational data and a power-law model from Insight-HXMT. Finally, Section 4 summarizes our conclusions.
2 The photon flux from PBH evaporation
@@
According to Hawking radiation theory, within the framework of the Standard Model, black holes emit all particles described by the Standard Model [13]. The particle emission spectrum resembles blackbody radiation, modulated by a greybody factor [14]:
$\frac{\partial^{2}N_{i}}{\partial E_{i}\partial t}=\frac{1}{2\pi}\frac{\Gamma_{i}\left(E_{i},M\right)}{e^{E_{i}/T_{H}}-(-1)^{2s}}.$ (2.1)
@@
The variables are defined as follows: ${i}$ is the type of particle, ${s}$ is the spin of the particle, ${N}$ is the number of particles, ${E}$ is the energy, ${\Gamma_{i}}$ is the greybody factor, ${M}$ is the mass of black hole, and ${T_{H}}$ is the Hawking temperature. The relationship between ${M}$ and ${T_{H}}$ is given by:
$T_{H}=\frac{M_{P}^{2}}{8\pi M},$ (2.2)
@@
Among the particles produced by Hawking radiation, some can give rise to unstable particles, such as hadrons formed through the fragmentation of gluons and quarks [15, 16]. These unstable particles undergo decay, producing stable secondary particles. In our work, we primarily focus on photons generated by Hawking radiation, including those directly emitted, secondary photons resulting from the decay of unstable particles produced by Hawking radiation, and photons produced via the annihilation of positrons emitted by Hawking radiation with electrons in the surrounding environment. The code BlackHawk employs to compute the photon and positron spectra generated by Hawking radiation [17].
2.1 Photons generated by Hawking radiation
@@
$\frac{\mathrm{d}\phi_{\gamma}^{\text{ext }}}{\mathrm{d}E}=\frac{f_{\mathrm{PBH}}\Omega_{\mathrm{DM}}\rho_{c}}{4\pi M}\int_{0}^{z_{\max}}\frac{\mathrm{d}z}{H(z)}\frac{\mathrm{d}^{2}N_{\gamma}}{\mathrm{d}E\mathrm{~{}d}t}(E(1+z)),$ (2.3)
where $f_{\mathrm{PBH}}$ denotes PBH abundance, representing the fraction of dark matter constituted by PBHs, and $N_{\gamma}$ is the total number of photons emitted per PBH. The Hubble parameter is expressed as $H(z)=H_{0}\sqrt{\Omega_{\Lambda}+\Omega_{M}(1+z)^{3}}$, with $H_{0}=67.36~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$, $\Omega_{\mathrm{DM}}=0.2645$, $\Omega_{\Lambda}=0.6847$ and $\rho_{c}=9.1\times 10^{-30}~{}\mathrm{g~{}{cm^{-3}}}$ [18]. The maximum redshift is chosen as $z_{\mathrm{max}}\geq 10$. Theoretically, $z_{\mathrm{max}}$ should correspond to the redshift at photon decoupling; however, this integral is not significantly sensitive to large values of $z_{\mathrm{max}}$.
@@
$\frac{\mathrm{d}\phi_{\gamma}^{\text{gal}}}{\mathrm{d}E}=\frac{f_{\mathrm{PBH}}\bar{J}_{D}}{{4\pi}M}\frac{\mathrm{d}^{2}N_{\gamma}}{\mathrm{d}E_{\gamma}\mathrm{d}t},$ (2.4)
where $\bar{J}_{D}$ is referred to as the astrophysical $J$-factor in some studies related to dark matter searches [19, 20, 21], typically used to describe the dark matter density per solid angle along the line of sight. Its specific form is :
$\bar{J}_{D}=\frac{1}{\Delta\Omega}\int_{\Delta\Omega}\mathrm{d}\Omega\int_{\text{l.o.s}}\mathrm{d}l\cdot\rho_{\mathrm{DM}}(l,\psi).$ (2.5)
Here, $l$ is the distance along the line of sight, typically integrated to infinity, and $\psi$ is the angle between the line of sight and the direction of the Galactic center. In our analysis, we adopt a Navarro-Frenk-White (NFW) profile [22] for the dark matter distribution, parameterized as:
$\rho_{\mathrm{DM}}(r)=\frac{\rho_{o}}{\frac{r}{r_{s}}\left(1+\frac{r}{r_{s}}\right)^{2}},$ (2.6)
where $\rho_{o}=0.839~{}\mathrm{GeV~{}cm^{-3}}$ is the characteristic density, and $r_{s}=11~{}\mathrm{kpc}$ is the scale radius [23].
It is important to note that, although the photon fluxes resulting from the evaporation of Galactic and extragalactic PBHs are theoretically expected to exhibit identical spectral features, a crucial distinction remains: ${\mathrm{d}\phi_{\gamma}^{\text{ext}}}/{\mathrm{d}E}$ represents the extragalactic photon flux integrated over redshifts from $z=0$ to $z=z_{\mathrm{max}}$, whereas ${\mathrm{d}\phi_{\gamma}^{\text{gal}}}/{\mathrm{d}E}$ corresponds solely to the Galactic photon flux at $z=0$. This difference leads to noticeably distinct spectral shapes, as shown in figure 1.
2.2 Photons from positron annihilations
When a positron and an electron interact, they can either annihilate directly, producing two 511 keV photons, or form a bound state known as positronium (Ps) before annihilation. If the spins of the electron and positron in the bound state are antiparallel, they form para-positronium, which annihilates into two photons with energies of 511 keV each. Conversely, if the spins are parallel, they form ortho-positronium, which annihilates into three photons with a broader energy distribution. According to ref. [24], when positrons produced by PBH evaporation interact with electrons in the cosmic medium, positronium formation ($f_{\mathrm{P}}=1$) is significantly more probable than direct annihilation ($f_{\mathrm{P}}=0$) under Milky Way-like conditons. Consequently, we consider only the scenario in which positronium is formed.
@@ 2.2 Photons from positron annihilations
$\frac{\mathrm{d}\phi_{\text{ann }}^{\text{ext }}}{\mathrm{d}E}={}$ $\frac{f_{\mathrm{PBH}}\Omega_{\mathrm{DM}}\rho_{c}}{4\pi M}\frac{\mathrm{d}N_{e^{+}}}{\mathrm{d}t}\int_{0}^{\mathrm{z}_{\max}}\frac{\mathrm{d}z}{H(z)}$ (2.7)
@@
$\frac{\mathrm{d}\phi_{\text{ann }}^{\text{gal}}}{\mathrm{d}E}=\frac{f_{\mathrm{PBH}}\bar{J}_{D}}{4\pi M}\frac{\mathrm{d}N_{e^{+}}}{\mathrm{d}t}\left(\frac{1}{2}\delta\left(E-m_{e}\right)+\frac{9}{4}\frac{h_{3\gamma}\left(E/{m_{e}}\right)}{m_{e}}\right).$ (2.8)
Here, $N_{e^{+}}$ represents the total number of positrons emitted by a single PBH, and $h_{3\gamma}(E/{m_{e}})$ denotes the energy distribution spectrum of photons produced by ortho-positronium annihilation [25, 26]. Its explicit form is given by
$h_{3\gamma}(x)$ $=\frac{2}{\pi^{2}-9}\Bigg{[}\frac{2-x}{x}+\frac{(1-x)x}{(2-x)^{2}}$ (2.9)
@@
$\psi^{\text{PBH}}=\frac{\mathrm{d}\phi^{\text{PBH}}}{\mathrm{d}E}=\frac{\mathrm{d}\phi_{\gamma}^{\text{ext }}}{\mathrm{d}E}+\frac{\mathrm{d}\phi_{\gamma}^{\text{gal}}}{\mathrm{d}E}+\frac{\mathrm{d}\phi_{\text{ann }}^{\text{ext }}}{\mathrm{d}E}+\frac{\mathrm{d}\phi_{\text{ann }}^{\text{gal}}}{\mathrm{d}E}.$ (2.10)
In figure 1, we plot the photon flux from various components with $M=10^{17}$ g and $f_{\mathrm{PBH}}=1$.
Figure 1: Various components for the total photon flux in Eq. (2.10) for $M=10^{17}$ g and $f_{\mathrm{PBH}}=1$. The red line represents the photon flux from Galactic sources, the blue line represents the photon flux from extragalactic sources. The solid line denotes the photon flux directly emitted by Hawking radiation, while the dashed line corresponds to the photon flux produced from positron annihilation.
From the figure, we observe that the photon flux directly emitted by Galactic PBHs exceeds that from extragalactic PBHs by more than an order of magnitude at the peak. Furthermore, for the total photon flux below 511 keV, the contribution from positron annihilation associated with Galactic and extragalactic PBHs dominates. According to Eqs. (2.7) and (2.8), the positron annihilation contributions depend solely on the electron spectrum generated by Hawking radiation. Consequently, whether the primary or secondary photon spectrum is used, the resulting constraints on $f_{\mathrm{PBH}}$ are expected to remain essentially unchanged.
3 Constrains with Insight-HXMT data

### v3

# thanks: Corresponding author

@@ Abstract
Primordial black holes (PBHs), a major candidate for dark matter, have been extensively constrained across most mass ranges. However, PBHs in the mass range $10^{17}$$10^{21}$ g remain a viable explanation for all dark matter. In this study, we use observational data from the Hard X-ray Modulation Telescope (Insight-HXMT) to refine constraints on PBHs within the mass range $2\times 10^{16}$$5\times 10^{17}$ g. Our analysis explores three scenarios: directly using observational data, incorporating the astrophysical background model (ABM), and employing the power-law spectrum with an exponential cutoff. Our research results indicate that although Insight-HXMT does not have an advantage in the first two scenarios, when considering the power-law model, its exceptional sensitivity in the hard X-ray regime and sufficiently high upper energy limit significantly strengthen the constraints on PBHs with masses greater than $10^{17}$ g compared to previous limits. Furthermore, the exclusion limit for PBHs as dark matter has reached $4\times 10^{17}$ g, which is comparable to the current threshold.
I Introduction
@@
Dark matter remains one of the most profound mysteries in modern astrophysics and cosmology. Observations such as Galactic rotation curves, gravitational lensing, and anisotropies in the cosmic microwave background provide compelling indirect evidence for its existence. However, despite these clues, the fundamental nature of dark matter continues to elude direct detection. Current models suggest that dark matter constitutes approximately 85% of the total matter in the universe WMAP:2010sfg; Planck:2015fie, interacting with baryonic matter primarily through gravity and remaining invisible to electromagnetic observations. Over the past few decades, numerous candidates have been proposed, such as weakly interacting massive particles (WIMPs), sterile neutrinos, and axions. These candidates often require modifications to the Standard Model of particle physics, and none have been definitively detected to date.
@@
Amid these ongoing searches, primordial black holes (PBHs) have emerged as an intriguing alternative, potentially bypassing the need for such modifications. PBHs are theoretical black holes that formed in the early universe due to the collapse of high-density fluctuations shortly after the Big Bang Hawking:1971ei. Unlike stellar black holes, PBHs can span a wide range of masses, from the Planck scale to billions of solar masses. This broad mass spectrum and early formation make PBHs relevant not only to the dark matter problem but also to various astrophysical phenomena. For instance, PBHs could account for the unexpected massive early galaxies observed by the James Webb Space Telescope (JWST) Yuan:2023bvh; Liu:2022bvr, and they offer a plausible explanation for the origin of the stellar-mass binary black holes detected by the LIGO and Virgo collaborations Sasaki:2016jop; Ali-Haimoud:2017rtz.
@@
While many mass ranges for PBHs as the entirety of dark matter have been ruled out through observations like lensing, gravitational waves, and accretion, a relatively unconstrained mass window remains between $10^{17}$ g and $10^{21}$Carr:2020gox; Carr:2021bzv. This mass range is particularly notable because it circumvents many of the astrophysical and cosmological constraints that apply to PBHs at both lower and higher masses. Although several previous studies have attempted to impose observational constraints on the abundance of PBHs within this mass window Laha:2019ssq; Laha:2020ivk; Coogan:2020tuf; Iguaz:2021irx; Berteaud:2022tws, the results have been unsatisfactory, yielding only marginal improvements. In this regime, Hawking radiation becomes a key mechanism for detecting PBHs and constraining their abundance. PBHs in this window are expected to emit radiation primarily in the keV–MeV energy range, making instruments sensitive to hard X-rays and soft $\gamma$-rays ideal for imposing tighter constraints on their abundance.
@@
The first Chinese X-ray astronomy satellite, the Hard X-ray Modulation Telescope (Insight-HXMT) offers a broad energy range, large effective area, and high sensitivity, particularly in the hard X-ray band Insight-HXMTTeam:2019dqg; Zhang:2019ikm. Its sensitivity surpasses that of other instruments, such as the International Gamma-Ray Astrophysics Laboratory/Imager on Board the INTEGRAL Satellite (INTEGRAL/IBIS) and Rossi X-ray Timing Explorer/High-Energy X-ray Timing Experiment (RXTE/HEXTE) Zhang:2019ikm. These features make Insight-HXMT particularly well-suited for detecting Hawking radiation from PBHs in the unconstrained mass window, positioning it as an ideal tool for constraining PBH abundance within this range.
@@
In this study, we use diffuse X-ray background data to constrain the abundance of PBHs in the mass range of $2\times 10^{16}$ g to $5\times 10^{17}$ g. Our analysis takes into account both the photon flux generated directly by Hawking radiation and the photons produced from the annihilation of positrons emitted during PBH evaporation. We include contributions to the diffuse X-ray background from both Galactic and extragalactic sources. The data used in our analysis are sourced from the diffuse X-ray background observed by Insight-HXMT Huang:2022guk. We evaluate three scenarios: directly using the observational data, incorporating an astrophysical background model (ABM), and applying the power-law model fitted from the data to extrapolate the upper energy limit. By leveraging Insight-HXMT’s advanced observational capabilities, our goal is to significantly improve existing constraints on PBH abundance and advance the limits set by prior studies.
@@
The structure of the paper is as follows: In Section II, we detail the calculations of photon flux from PBH evaporation. Section III presents the statistical analysis and constraints on PBH abundance using observational data and a power-law model from Insight-HXMT. Finally, Section IV summarizes our conclusions.
II The photon flux from PBH evaporation
@@
According to Hawking radiation theory, within the framework of the Standard Model, black holes emit all particles described by the Standard Model Hawking:1974rv. The particle emission spectrum resembles blackbody radiation, modulated by a greybody factor Hawking:1975vcx:
$\frac{\partial^{2}N_{i}}{\partial E_{i}\partial t}=\frac{1}{2\pi}\frac{\Gamma_{i}\left(E_{i},M\right)}{e^{E_{i}/T_{H}}-(-1)^{2s}}.$ (1)
@@
The variables are defined as follows: i{i} is the type of particle, s{s} is the spin of the particle, N{N} is the number of particles, E{E} is the energy, $\Gamma_{i}$} is the greybody factor, M{M} is the mass of the black hole, and $T_{H}$} is the Hawking temperature. The relationship between M{M} and $T_{H}$} is given by:
$T_{H}=\frac{M_{P}^{2}}{8\pi M},$ (2)
@@
Among the particles produced by Hawking radiation, some can give rise to unstable particles, such as hadrons formed through the fragmentation of gluons and quarks MacGibbon:1990zk; MacGibbon:1991tj. These unstable particles undergo decay, producing stable secondary particles. In our work, we primarily focus on photons generated by Hawking radiation, including those directly emitted, secondary photons resulting from the decay of unstable particles produced by Hawking radiation, and photons produced via the annihilation of positrons emitted by Hawking radiation with electrons in the surrounding environment. Photons from other processes, such as in-flight annihilation, are not included in this study as their contribution to the total photon flux is negligible. Detailed calculations in Ref. Tan:2024nbx show that the photon flux yield from in-flight annihilation is approximately four orders of magnitude lower than the primary photon flux of Hawking radiation. We use the code BlackHawk to compute the photon and positron spectra generated by Hawking radiation Arbey:2019mbc.
II.1 Photons generated by Hawking radiation
@@
$\frac{\mathrm{d}\phi_{\gamma}^{\text{ext }}}{\mathrm{d}E}=\frac{f_{\mathrm{PBH}}\Omega_{\mathrm{DM}}\rho_{c}}{4\pi M}\int_{0}^{z_{\max}}\frac{\mathrm{d}z}{H(z)}\frac{\partial^{2}N_{\gamma}}{\partial E\partial t}(E(1+z)),$ (3)
where $f_{\mathrm{PBH}}$ denotes PBH abundance, representing the fraction of dark matter constituted by PBHs, and $N_{\gamma}$ is the total number of photons emitted per PBH. The Hubble parameter is $H(z)=H_{0}\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}$, with $H_{0}=67.36~\mathrm{km\,s^{-1}\,Mpc^{-1}}$ and $\Omega_{\Lambda}=0.6847$. The matter density parameter is $\Omega_{m}=0.315$, which is the sum of $\Omega_{\mathrm{DM}}=0.2645$ for dark matter and $\Omega_{b}\simeq 0.0505$ for baryons. The critical density is $\rho_{c}=9.1\times 10^{-30}~\mathrm{g\,cm^{-3}}$ Planck:2018vyg.
The maximum redshift is chosen as $z_{\mathrm{max}}\geq 10$. Theoretically, $z_{\mathrm{max}}$ should correspond to the redshift at photon decoupling; however, this integral is not significantly sensitive to large values of $z_{\mathrm{max}}$.
@@
$\frac{\mathrm{d}\phi_{\gamma}^{\text{gal}}}{\mathrm{d}E}=\frac{f_{\mathrm{PBH}}\bar{J}_{D}}{{4\pi}M}\frac{\partial^{2}N_{\gamma}}{\partial E_{\gamma}\partial t},$ (4)
where $\bar{J}_{D}$ is referred to as the astrophysical $J$-factor in some studies related to dark matter searches Coogan:2021sjs; Boddy:2015efa; CTAO:2024wvb, typically used to describe the dark matter density per solid angle along the line of sight. Its specific form is:
$\bar{J}_{D}=\frac{1}{\Delta\Omega}\int_{\Delta\Omega}\mathrm{d}\Omega\int_{\text{l.o.s.}}\mathrm{d}l\,\rho_{\mathrm{DM}}(l,\psi).$ (5)
Here, $l$ is the distance along the line of sight, typically integrated to infinity, and $\psi$ is the angle between the line of sight and the direction of the Galactic center. In our analysis, we adopt a Navarro-Frenk-White (NFW) profile Navarro:1996gj for the dark matter distribution, parameterized as:
$\rho_{\mathrm{DM}}(r)=\frac{\rho_{0}}{\frac{r}{r_{s}}\left(1+\frac{r}{r_{s}}\right)^{2}},$ (6)
where $\rho_{0}=0.839~\mathrm{GeV\,cm^{-3}}$ is the characteristic density, and $r_{s}=11~\mathrm{kpc}$ is the scale radius deSalas:2019pee.
It is important to note that, although the photon fluxes resulting from the evaporation of Galactic and extragalactic PBHs are theoretically expected to exhibit identical spectral features, a crucial distinction remains: ${\mathrm{d}\phi_{\gamma}^{\text{ext}}}/{\mathrm{d}E}$ represents the extragalactic photon flux integrated over redshifts from $z=0$ to $z=z_{\mathrm{max}}$, whereas ${\mathrm{d}\phi_{\gamma}^{\text{gal}}}/{\mathrm{d}E}$ corresponds solely to the Galactic photon flux at $z=0$. This difference leads to noticeably distinct spectral shapes, as shown in Fig. 1.
II.2 Photons from positron annihilations
When a positron and an electron interact, they can either annihilate directly, producing two 511 keV photons, or form a bound state known as positronium (Ps) before annihilation. If the spins of the electron and positron in the bound state are antiparallel, they form para-positronium, which annihilates into two photons with energies of 511 keV each. Conversely, if the spins are parallel, they form ortho-positronium, which annihilates into three photons with a broader energy distribution. According to Ref. Guessoum:2005cb, when positrons produced by PBH evaporation interact with electrons in the cosmic medium, positronium formation ($f_{\mathrm{P}}=1$) is significantly more probable than direct annihilation ($f_{\mathrm{P}}=0$) under Milky Way-like conditions. Consequently, we consider only the scenario in which positronium is formed.
@@ II.2 Photons from positron annihilations
$\frac{\mathrm{d}\phi_{\text{ann }}^{\text{ext }}}{\mathrm{d}E}=$ $\frac{f_{\mathrm{PBH}}\Omega_{\mathrm{DM}}\rho_{c}}{4\pi M}\frac{\mathrm{d}N_{e^{+}}}{\mathrm{d}t}\int_{0}^{\mathrm{z}_{\max}}\frac{\mathrm{d}z}{H(z)}$ (7)
@@
$\frac{\mathrm{d}\phi_{\text{ann }}^{\text{gal}}}{\mathrm{d}E}=\frac{f_{\mathrm{PBH}}\bar{J}_{D}}{4\pi M}\frac{\mathrm{d}N_{e^{+}}}{\mathrm{d}t}\left(\frac{1}{2}\delta\left(E-m_{e}\right)+\frac{9}{4}\frac{h_{3\gamma}\left(E/{m_{e}}\right)}{m_{e}}\right).$ (8)
Here, $N_{e^{+}}$ represents the total number of positrons emitted by a single PBH, and $h_{3\gamma}(E/{m_{e}})$ denotes the energy distribution spectrum of photons produced by ortho-positronium annihilation Ore:1949te; Manohar:2003xv. Its explicit form is given by
$h_{3\gamma}(x)$ $=\frac{2}{\pi^{2}-9}\Bigg[\frac{2-x}{x}+\frac{(1-x)x}{(2-x)^{2}}$ (9)
@@
$\psi^{\text{PBH}}=\frac{\mathrm{d}\phi^{\text{PBH}}}{\mathrm{d}E}=\frac{\mathrm{d}\phi_{\gamma}^{\text{ext }}}{\mathrm{d}E}+\frac{\mathrm{d}\phi_{\gamma}^{\text{gal}}}{\mathrm{d}E}+\frac{\mathrm{d}\phi_{\text{ann }}^{\text{ext }}}{\mathrm{d}E}+\frac{\mathrm{d}\phi_{\text{ann }}^{\text{gal}}}{\mathrm{d}E}.$ (10)
In Fig. 1, we plot the photon flux from various components with $M=10^{17}$ g and $f_{\mathrm{PBH}}=1$.
Figure 1: Various components of the total photon flux in Eq. (10) for $M=10^{17}$ g and $f_{\mathrm{PBH}}=1$. The red line represents the photon flux from Galactic sources, the blue line represents the photon flux from extragalactic sources. The solid line denotes the photon flux directly emitted by Hawking radiation, while the dashed line corresponds to the photon flux produced from positron annihilation.
From the figure, we observe that the photon flux directly emitted by Galactic PBHs exceeds that from extragalactic PBHs by more than an order of magnitude at the peak. Furthermore, for the total photon flux below 511 keV, the contribution from positron annihilation associated with Galactic and extragalactic PBHs dominates. In the Insight-HXMT energy band (1–250 keV), the sensitivity is dominated by positron–electron annihilation, and the resulting 95% CL limits are primarily set by the annihilation component.
III Constraints with Insight-HXMT data